OpenStudy (anonymous):

Find a polynomial p of degree 3 such that -3, -1 and 7 are zeros of p and p(1)=2

OpenStudy (anonymous):

p(1)-2=0

OpenStudy (anonymous):

zoom

OpenStudy (amistre64):

(-3+3)(-1+1)(7-7) r r r (x+3)(x+1)(x-7) 1 1 1 --------------- 4 2 6 = 48 so divide by 24

OpenStudy (amistre64):

1/24 (x+3)(x+1)(x-7)

OpenStudy (amistre64):

not subtractoin; scaled

OpenStudy (anonymous):

\[a(x+3)(x+1)(x-7)\] solve \[p(1)=2\] for a \[p(1)=a\times 4\times 2\times -6=2\] etecs

OpenStudy (anonymous):

actually i get \[a=-\frac{1}{24}\]

OpenStudy (amistre64):

meh ... lol

OpenStudy (anonymous):

on account of 1 - 7 = -6 but what is a minus sign between friends?

OpenStudy (amistre64):

thats what we got lawyers for :)

OpenStudy (anonymous):

i knew they were there for some reason...

OpenStudy (anonymous):

so what do I do after this step? 1/24 (x+3)(x+1)(x-7)

OpenStudy (anonymous):

(x+1) (x+3)(x-7) = (x+1) ( x^2 - 4x - 21) = x^3 - 4x^2 - 21x + ( x^2 - 4x - 21) = x^3 - 3x^2 - 25x - 21